The width of a rectangle is 6 less than twice its length. If the area of the rectangle is 24 cm^2, what is the length of the diagonal?
Answer by stanbon(75887)

The width of a rectangle is 6 less than twice its length. If the area of the rectangle is 24 cm^2, what is the length of the diagonal?
--------------
length = "x"
width = "2x-6"
-----
Equation:
x(2x-6) = 24
-----
2x^2 - 6x - 24 = 0
x^2-3x-12 = 0
---------
x = [3+-sqrt(9-4*-12)]/2
---------
x = length = [3+-sqrt(57)]/2
-----
width = 2[length]-6
-----------------
diagonal = sqrt[length^2 + width^2]
Cheers,
Stan H.
-------------